### v1

@@ V-B LLM for Telecom Security and Attack Detection
Recently, NLP has achieved numerous successes in capturing informative features from multi-modal and heterogeneous data across various application scenarios, including sentiment analysis, speech recognition, and machine translation, among others. Specifically, LLM techniques have emerged as a promising direction across various NLP applications, which are beneficial to explaining textual inputs and transforming them into quantitative representations. The common method to apply LLM models across various domains involves employing general-domain models as baselines, followed by fine-tuning them for specific domain tasks. To enhance the security of telecom networks through LLM techniques, it is important to note that the security language, such as ransomware, API, OAuth, and keylogger, significantly differs in structure and semantics from the general linguistic language. This suggests that a conventional LLM may find it challenging to fully understand the specific vocabulary inherent to security-related texts, potentially leading to limited generalization ability in security applications. To this end, existing studies that employ LLM techniques for attack detection can be categorized into two primary directions as follows:
V-B1 Fine-tuning pre-trained LLM models
@@
V-B2 Building security-specific LLM models from scratch
In addition to fine-tuning, another strategy is to build an LLM from scratch specifically designed for network-based attack detection. For example, Ferrag et al. [124] designed SecurityBERT for detecting the ever-evolving cyber threat landscape, which involves several steps: data preparation, data tokenization, model training, and model deployment, as shown in Fig. 7. In particular, the authors utilize a publically available dataset EdgeIIoTset [140] related to the Internet of Things (IoT) and Industrial IoT (IIoT) connectivity protocols, categorized into five types of threats: DoS/D-DoS attacks, information gathering, MITM attacks, injection attacks, and malware attacks. Then, to leverage the power of LLM models, null features are eliminated during the feature extraction in [124], and both numerical and categorical features are converted into textual representations. Specifically, each feature is combined with its column name and value and then subjected to hashing. The hashed values from the same instance are merged into a sequence, which generates a fixed-length textual representation of the network traffic data while maintaining privacy. After that, ByteLevelBPETokenizer [141] is subsequently applied to segment the textual representations of the network traffic data. This segmentation process breaks down the text into smaller subwords, expected to be found in the tokenizer’s vocabulary. After the pre-training phase, the model is fine-tuned on a labelled dataset [140] by adding a Softmax activation function at the output layer, which allows SecurityBERT to enable the learned contextual representations in the specific task of attack detection. Finally, in the deployment phase of Fig. 7, once attacks are identified through SecurityBERT, FalconLLM is further employed to determine the severity and negative impact of identified attacks, leading to the formulation of potential mitigation strategies and recovery procedures.
SecurityBERT is employed to identify normal events and 14 distinct attacks in [124], such as DDoS_UDP, DDoS_ICMP, SQL_Injection, Vulnerability_Scanner, etc. The experiment shows that SecurityBERT achieves the average accuracy, recall, and F1-score of 0.98, 0.84, and 0.84, respectively, demonstrating the strong classification capabilities of security-specific LLM models in identifying various types of attacks. In addition, SecurityBERT significantly surpasses the performance of traditional ML and deep learning models such as decision tree, convolutional neural networks (CNN), recurrent neural network (RNN), and long short-term memory (LSTM).
@@ V-C1 Using LLMs for telecom user feedback classification and analyses
Understanding user feedback is crucial for telecom operators to improve the QoE and maintain customer satisfaction and loyalty. For instance, Vieira et al. applied CNN and LSTM networks in [127] for sentiment analysis and topic classification, and the analysis proved that 78.3% of the complaints are related to weak signal coverage, and 92% of these regions have coverage problems considering a specific cellular operator. These analyses can be particularly useful for telecom operators to improve service quality such as signal coverage and strength. However, user feedback can be complicated by involving service experiences, suggestions, recommendations, and complaints. In addition, the feedback can be collected from various sources, e.g., social media, websites, phone calls, and company collection. These challenges require more advanced ML techniques to better classify and capture user’s intentions. The LLM has shown superb performance in a range of text-related tasks, e.g., question answering, summarization, dialogue, and sentiment analysis, outperforming many existing techniques even in zero-shot settings. For instance, Aftan et. al applied AraBERT model to classify telecom customer satisfaction in Saudi Arabia by using the Twitter dataset [126], and the BERT-based model obtained more accurate and stable results than conventional CNN and RNN algorithms. In addition, using LLM models to analyze customers’ experience and intent has attracted considerable interest from both industry and academia, e.g., Microsoft has proposed to use LLMs to generate, validate, and apply user intent taxonomies[145]. Therefore, it shows great promise in integrating LLM technologies into the telecom industry for text-related classification tasks.
@@ V-C2 LLM-aided telecom standard classification
Telecom standards refer to agreed-upon specifications that ensure the interoperability, security, and reliability of telecom services. Standards play a critical role in global telecoms [111], such as 3G, 4G, and emerging 5G for mobile communications, IEEE 802.11 for Wi-Fi, and ITU-T recommendations. For instance, 3GPP is the main organization for telecom standard development, which includes three technical specification groups, and each specification group consists of multiple working groups. Given the large number of existing specifications with diverse topics, Lina et al. proposed to use LLM models for specification classification, classifying the text into an existing working group automatically[13]. Fig.8 summarized the key processes of using LLM models to classify the 3GPP specifications. With proper pre-processing and fine-tuning, the experiment in [13] can achieve an 80% accuracy even if only 20% text segments are used. The experiment results also prove that increasing the length of technical text segments can significantly improve classification accuracy.
Textual descriptions and documents are frequently involved in the telecom industry, e.g., user comments, standard specifications, technical and troubleshooting reports, etc. Incorporating LLM models for text processing and classification will contribute to more intelligent and reliable telecom networks.
@@ V-D Image Classification
The wireless environment can be very complicated with walking pedestrians, moving vehicles, building blockages, and other obstacles. Therefore, it requires dedicated model training and fine-tuning to extract useful information and identify specific objects. LLM models have been pre-trained on a huge amount of real-world datasets, and some LLMs, such as flamingo [136] and GPT-4V [149], have proved versatile capabilities on various vision-related tasks, e.g., using text to generate images, describing given images, and object detection. For instance, Matsuura et al. investigate the zero-shot image classification capabilities of the LLaVA model [26], and they found that the performance can be significantly improved with a combination of carefully crafted prompts, hierarchical classification strategies, and adjusted model temperatures. Meanwhile, Pratt et al. [129] also demonstrate that using LLM’s knowledge can immediately improve zero-shot accuracy on a variety of image classification tasks, saving considerable manual effort. In addition, LLMs can also describe and summarize the image content for further classification, documentation, and processing, and an example is given in [17] by generating the accident report of a car crash.
Finally, Fig.9 presents an example of using LLM models for image classification and object detection in radio access networks. In particular, the cameras attached to the BS can capture environmental images, and then the image data will be sent to the LLM at the network edge by wired backhaul. The LLM can use computational resources at the edge cloud for image processing, classification, and detecting object locations such as vehicles, users, and blockage buildings. After that, the edge cloud can send back the classification and detection results to BSs, and then the BS can adjust the beamforming and hand-off decisions accordingly.
@@ V-E Encrypted Traffic Classification
Attack detection Attack detection is vital for maintaining the security and reliability of telecom networks. LLMs have showcased a strong ability to capture discriminative information within multi-modal and heterogeneous network data. Fine-tuning LLMs on security-specific language emerges as a promising approach for attack detection. This method allows fine-tuned LLM models to maintain their proficiency in processing general English vocabulary while excelling at achieving specific security objectives. LLM techniques have the strong advantages of processing both numerical traffic loads and descriptive security-related textual contents, e.g., ransomware and keylogger, achieving better performance than existing ML and DL algorithms. LLM models can be effectively employed for detecting cyber attacks [121, 124] and contributing to the mitigation and recovery strategies against such attacks [124].
@@
Text classification Text classification and processing are very useful for the telecom industry. LLM models have shown great promise in understanding and processing text and languages. Fine-tuning LLM models on telecom language is required, e.g., network trouble report datasets and 3GPP technical specifications. There are no specific requirements for prompts. Automatic text classification and processing will greatly save human efforts on many document-related tasks, e.g., automated troubleshooting report generation and ranking. The telecom applications include user enquiries and intent classification and analyses[127], automated trouble report classification[144], standard specification classification[30].
@@
Image classification Computer vision is a very useful technique for 6G networks, enabling 3D sensing for the environment. Some LLMs have shown impressive capabilities in image and vision-related tasks. The study in [26] shows that carefully crafted prompts are critical to improving the classification performance of LLM models, e.g., ”Fill in the blank, this is a picture of {} ”. However, fine-tuning LLM models on telecom-image datasets can improve classification accuracy. LLM’s zero-shot learning capability can avoid the complexity of dedicated model training. For instance, [26] achieved a satisfactory performance by pure prompt engineering without any fine-tuning. LLM-aided image classification can be used for blockage prediction[146], proactive beamforming and hand-off[25], user localization [128], etc.
@@
Traffic classification Network traffic classification is an essential technique in network management and security, which aims at identifying the category of traffic from various applications. LLM techniques have demonstrated remarkable performance in encrypted traffic classification. Fine-tuning LLM models on labeled network data is crucial for ensuring their adaptability across various traffic classification scenarios, such as single packet and single flow classification. LLMs are capable of learning generic traffic representations from extensive amounts of unlabeled, encrypted traffic without plaintext, resulting in extracting valuable insights from encrypted traffic for downstream traffic classification. Traffic analyses and classification are very common tasks in telecom networks. LLMs can be effectively applied for encrypted traffic classification [131, 130].
@@ V-F Discussions and Analyses
In particular, Section V-B demonstrates that LLM techniques have great potential for telecom network security. Security is an important topic for telecom operations, and LLM models can contribute through their classification and detection capabilities. In particular, the LLM can handle multi-modal and heterogeneous network data, e.g., CSI, traffic load level, network device logs and network configurations, and then extract useful network security information from these correlated inputs. Additionally, some LLMs can also recommend response and recovery strategies for network incidents[124]. This indicates the potential of building an end-to-end telecom security system, from status monitoring and attack detection to incident response and recovery.
Meanwhile, LLM can serve as a zero-shot classifier. Telecom networks indicate a complicated dynamic environment, leading to various tasks. Existing methods are usually task-specific, with dedicated designs for each incoming request. By contrast, some LLM models have shown zero-shot classification capabilities[26]. For instance, they can be directly used to classify images captured by the cameras on the BS, or analyze customer comments without prior training. Such a feature can be very useful in handling diverse tasks in complicated telecom systems such as object detection and user localization. In addition, LLM models have outstanding capabilities in processing text-related tasks, including both natural languages, such as customer comments and system language like network log files. These textual tasks are usually performed manually, but LLM models can easily handle different classification and detection tasks with much less human intervention.
Finally, LLM models can contribute to vision-aided telecom. Sensing is increasingly important for wireless networks, and computer vision is an important approach to capturing wireless environment dynamics. With pre-trained real-world knowledge, LLM models can be directly used for image and vision-related tasks, such as image description, image-text transformation, object detection, and image classification. In addition, LLM technologies also have advantages over conventional algorithms in terms of generalization capabilities. This means that LLMs can process various telecom tasks without extra training, e.g., blockage detection and prediction by using BS cameras[146], proactive beamforming and hand-off[25], and user localization [128].
@@ VI-A Motivations and Optimization Capabilities of LLM
[159] It proposed a novel framework to reinforce language agents through linguistic feedback. The agent verbally reflects on task feedback signals, maintaining the reflective text in an episodic memory buffer to induce better decision-making. The proposed framework achieves a 91% accuracy on the HumanEval coding benchmark, surpassing the previous state-of-the-art GPT-4 that achieves 80%. LLM models have self-improvement capability, which means they can work as an agent to receive network environment feedback and improve the policies based on textual input.
@@
[162] Evaluating the optimization capabilities of LLM models across diverse tasks and data sizes, including gradient descent, hill-climbing, grid-search, and black-box optimization. 1) The LLM show strong optimization capabilities; 2) LLM models perform well in small-size samples; 3) They exhibit strong performance in gradient-descent; 4) LLMs are black-box optimizers. Black-box optimizer is a promising approach to estimating the unknown loss function, which is especially useful since telecom networks become more complicated.
@@
Secondly, LLMs can lower the training and fine-tuning difficulties of ML-based network optimization. Algorithm training is considered one of the main obstacles to realizing AI-enabled wireless networks, which is usually very time-consuming. By contrast, LLM has shown impressive few-shot or even zero-shot learning capabilities in many fields[167]. In particular, LLM models can learn in context from few or zero network management and optimization examples and then generalize to incoming new tasks. By providing a handful of examples, the LLM agent can quickly learn the hidden patterns without any extra model training and fine-tuning, saving considerable time and effort for algorithm training in network management. In addition, such fast learning capability is critical to make rapid responses to network dynamics. This means that network optimization decisions can be efficiently adjusted based on traffic patterns, user types, and operator demands.

### v2

@@ V-B LLM for Telecom Security and Attack Detection
Recently, NLP has achieved numerous successes in capturing informative features from multi-modal and heterogeneous data across various application scenarios, including sentiment analysis, speech recognition, and machine translation, among others. Specifically, LLM techniques have emerged as a promising direction across various NLP applications, which are beneficial to explaining textual inputs and transforming them into quantitative representations. The common method to apply LLMs across various domains involves employing general-domain models as baselines, followed by fine-tuning them for specific domain tasks. To enhance the security of telecom networks through LLM techniques, it is important to note that the security language, such as ransomware, API, OAuth, and keylogger, significantly differs in structure and semantics from the general linguistic language. This suggests that a conventional LLM may find it challenging to fully understand the specific vocabulary inherent to security-related texts, potentially leading to limited generalization ability in security applications. To this end, existing studies that employ LLM techniques for attack detection can be categorized into two primary directions as follows:
V-B1 Fine-tuning pre-trained LLMs
@@
V-B2 Building security-specific LLMs from scratch
In addition to fine-tuning, another strategy is to build an LLM from scratch specifically designed for network-based attack detection. For example, Ferrag et al. [139] designed SecurityBERT for detecting the ever-evolving cyber threat landscape, which involves several steps: data preparation, data tokenization, model training, and model deployment, as shown in Fig. 8. In particular, the authors utilize a publically available dataset EdgeIIoTset [155] related to the Internet of Things (IoT) and Industrial IoT (IIoT) connectivity protocols, categorized into five types of threats: DoS/D-DoS attacks, information gathering, MITM attacks, injection attacks, and malware attacks. Then, to leverage the power of LLMs, null features are eliminated during the feature extraction in [139], and both numerical and categorical features are converted into textual representations. Specifically, each feature is combined with its column name and value and then subjected to hashing. The hashed values from the same instance are merged into a sequence, which generates a fixed-length textual representation of the network traffic data while maintaining privacy. After that, ByteLevelBPETokenizer [156] is subsequently applied to segment the textual representations of the network traffic data. This segmentation process breaks down the text into smaller subwords, expected to be found in the tokenizer’s vocabulary. After the pre-training phase, the model is fine-tuned on a labelled dataset [155] by adding a Softmax activation function at the output layer, which allows SecurityBERT to enable the learned contextual representations in the specific task of attack detection. Finally, in the deployment phase of Fig. 8, once attacks are identified through SecurityBERT, FalconLLM is further employed to determine the severity and negative impact of identified attacks, leading to the formulation of potential mitigation strategies and recovery procedures.
SecurityBERT is employed to identify normal events and 14 distinct attacks in [139], such as DDoS_UDP, DDoS_ICMP, SQL_Injection, Vulnerability_Scanner, etc. The experiment shows that SecurityBERT achieves the average accuracy, recall, and F1-score of 0.98, 0.84, and 0.84, respectively, demonstrating the strong classification capabilities of security-specific LLMs in identifying various types of attacks. In addition, SecurityBERT significantly surpasses the performance of traditional ML and deep learning models such as decision tree, convolutional neural networks (CNN), recurrent neural network (RNN), and long short-term memory (LSTM).
@@ V-C1 Using LLMs for telecom user feedback classification and analyses
Understanding user feedback is crucial for telecom operators to improve the QoE and maintain customer satisfaction and loyalty. For instance, Vieira et al. applied CNN and LSTM networks in [142] for sentiment analysis and topic classification, and the analysis proved that 78.3% of the complaints are related to weak signal coverage, and 92% of these regions have coverage problems considering a specific cellular operator. These analyses can be particularly useful for telecom operators to improve service quality such as signal coverage and strength. However, user feedback can be complicated by involving service experiences, suggestions, recommendations, and complaints. In addition, the feedback can be collected from various sources, e.g., social media, websites, phone calls, and company collection. These challenges require more advanced ML techniques to better classify and capture user’s intentions. The LLM has shown superb performance in a range of text-related tasks, e.g., question answering, summarization, dialogue, and sentiment analysis, outperforming many existing techniques even in zero-shot settings. For instance, Aftan et. al applied AraBERT model to classify telecom customer satisfaction in Saudi Arabia by using the Twitter dataset [141], and the BERT-based model obtained more accurate and stable results than conventional CNN and RNN algorithms. In addition, using LLMs to analyze customers’ experience and intent has attracted considerable interest from both industry and academia, e.g., Microsoft has proposed to use LLMs to generate, validate, and apply user intent taxonomies[160]. Therefore, it shows great promise in integrating LLM technologies into the telecom industry for text-related classification tasks.
@@ V-C2 LLM-aided telecom standard classification
Telecom standards refer to agreed-upon specifications that ensure the interoperability, security, and reliability of telecom services. Standards play a critical role in global telecoms [126], such as 3G, 4G, and emerging 5G for mobile communications, IEEE 802.11 for Wi-Fi, and ITU-T recommendations. For instance, 3GPP is the main organization for telecom standard development, which includes three technical specification groups, and each specification group consists of multiple working groups. Given the large number of existing specifications with diverse topics, Lina et al. proposed to use LLMs for specification classification, classifying the text into an existing working group automatically[13]. Fig.9 summarized the key processes of using LLMs to classify the 3GPP specifications. With proper pre-processing and fine-tuning, the experiment in [13] can achieve an 80% accuracy even if only 20% text segments are used. The experiment results also prove that increasing the length of technical text segments can significantly improve classification accuracy.
Textual descriptions and documents are frequently involved in the telecom industry, e.g., user comments, standard specifications, technical and troubleshooting reports, etc. Incorporating LLMs for text processing and classification will contribute to more intelligent and reliable telecom networks.
@@ V-D Image Classification
The wireless environment can be very complicated with walking pedestrians, moving vehicles, building blockages, and other obstacles. Therefore, it requires dedicated model training and fine-tuning to extract useful information and identify specific objects. LLMs have been pre-trained on a huge amount of real-world datasets, and some LLMs, such as flamingo [151] and GPT-4V [164], have proved versatile capabilities on various vision-related tasks, e.g., using text to generate images, describing given images, and object detection. For instance, Matsuura et al. investigate the zero-shot image classification capabilities of the LLaVA model [26], and they found that the performance can be significantly improved with a combination of carefully crafted prompts, hierarchical classification strategies, and adjusted model temperatures. Meanwhile, Pratt et al. [144] also demonstrate that using LLM’s knowledge can immediately improve zero-shot accuracy on a variety of image classification tasks, saving considerable manual effort. In addition, LLMs can also describe and summarize the image content for further classification, documentation, and processing, and an example is given in [17] by generating the accident report of a car crash.
Finally, Fig.10 presents an example of using LLMs for image classification and object detection in radio access networks. In particular, the cameras attached to the BS can capture environmental images, and then the image data will be sent to the LLM at the network edge by wired backhaul. The LLM can use computational resources at the edge cloud for image processing, classification, and detecting object locations such as vehicles, users, and blockage buildings. After that, the edge cloud can send back the classification and detection results to BSs, and then the BS can adjust the beamforming and hand-off decisions accordingly.
@@ V-E Encrypted Traffic Classification
Attack detection Attack detection is vital for maintaining the security and reliability of telecom networks. LLMs have showcased a strong ability to capture discriminative information within multi-modal and heterogeneous network data. Fine-tuning LLMs on security-specific language emerges as a promising approach for attack detection. This method allows fine-tuned LLMs to maintain their proficiency in processing general English vocabulary while excelling at achieving specific security objectives. LLM techniques have the strong advantages of processing both numerical traffic loads and descriptive security-related textual contents, e.g., ransomware and keylogger, achieving better performance than existing ML and DL algorithms. LLMs can be effectively employed for detecting cyber attacks [136, 139] and contributing to the mitigation and recovery strategies against such attacks [139].
@@
Text classification Text classification and processing are very useful for the telecom industry. LLMs have shown great promise in understanding and processing text and languages. Fine-tuning LLMs on telecom language is required, e.g., network trouble report datasets and 3GPP technical specifications. There are no specific requirements for prompts. Automatic text classification and processing will greatly save human efforts on many document-related tasks, e.g., automated troubleshooting report generation and ranking. The telecom applications include user enquiries and intent classification and analyses[142], automated trouble report classification[159], standard specification classification[30].
@@
Image classification Computer vision is a very useful technique for 6G networks, enabling 3D sensing for the environment. Some LLMs have shown impressive capabilities in image and vision-related tasks. The study in [26] shows that carefully crafted prompts are critical to improving the classification performance of LLMs, e.g., ”Fill in the blank, this is a picture of {} ”. However, fine-tuning LLMs on telecom-image datasets can improve classification accuracy. LLM’s zero-shot learning capability can avoid the complexity of dedicated model training. For instance, [26] achieved a satisfactory performance by pure prompt engineering without any fine-tuning. LLM-aided image classification can be used for blockage prediction[161], proactive beamforming and hand-off[25], user localization [143], etc.
@@
Traffic classification Network traffic classification is an essential technique in network management and security, which aims at identifying the category of traffic from various applications. LLM techniques have demonstrated remarkable performance in encrypted traffic classification. Fine-tuning LLMs on labeled network data is crucial for ensuring their adaptability across various traffic classification scenarios, such as single packet and single flow classification. LLMs are capable of learning generic traffic representations from extensive amounts of unlabeled, encrypted traffic without plaintext, resulting in extracting valuable insights from encrypted traffic for downstream traffic classification. Traffic analyses and classification are very common tasks in telecom networks. LLMs can be effectively applied for encrypted traffic classification [146, 145].
@@ V-F Discussions and Analyses
In particular, Section V-B demonstrates that LLM techniques have great potential for telecom network security. Security is an important topic for telecom operations, and LLMs can contribute through their classification and detection capabilities. In particular, the LLM can handle multi-modal and heterogeneous network data, e.g., CSI, traffic load level, network device logs and network configurations, and then extract useful network security information from these correlated inputs. Additionally, some LLMs can also recommend response and recovery strategies for network incidents[139]. This indicates the potential of building an end-to-end telecom security system, from status monitoring and attack detection to incident response and recovery.
Meanwhile, LLM can serve as a zero-shot classifier. Telecom networks indicate a complicated dynamic environment, leading to various tasks. Existing methods are usually task-specific, with dedicated designs for each incoming request. By contrast, some LLMs have shown zero-shot classification capabilities[26]. For instance, they can be directly used to classify images captured by the cameras on the BS, or analyze customer comments without prior training. Such a feature can be very useful in handling diverse tasks in complicated telecom systems such as object detection and user localization. In addition, LLMs have outstanding capabilities in processing text-related tasks, including both natural languages, such as customer comments and system language like network log files. These textual tasks are usually performed manually, but LLMs can easily handle different classification and detection tasks with much less human intervention.
Finally, LLMs can contribute to vision-aided telecom. Sensing is increasingly important for wireless networks, and computer vision is an important approach to capturing wireless environment dynamics. With pre-trained real-world knowledge, LLMs can be directly used for image and vision-related tasks, such as image description, image-text transformation, object detection, and image classification. In addition, LLM technologies also have advantages over conventional algorithms in terms of generalization capabilities. This means that LLMs can process various telecom tasks without extra training, e.g., blockage detection and prediction by using BS cameras[161], proactive beamforming and hand-off[25], and user localization [143].
@@ VI-A Motivations and Optimization Capabilities of LLM
[174] It proposed a novel framework to reinforce language agents through linguistic feedback. The agent verbally reflects on task feedback signals, maintaining the reflective text in an episodic memory buffer to induce better decision-making. The proposed framework achieves a 91% accuracy on the HumanEval coding benchmark, surpassing the previous state-of-the-art GPT-4 that achieves 80%. LLMs have self-improvement capability, which means they can work as an agent to receive network environment feedback and improve the policies based on textual input.
@@
[177] Evaluating the optimization capabilities of LLMs across diverse tasks and data sizes, including gradient descent, hill-climbing, grid-search, and black-box optimization. 1) The LLM show strong optimization capabilities; 2) LLMs perform well in small-size samples; 3) They exhibit strong performance in gradient-descent; 4) LLMs are black-box optimizers. Black-box optimizer is a promising approach to estimating the unknown loss function, which is especially useful since telecom networks become more complicated.
@@
Secondly, LLMs can lower the training and fine-tuning difficulties of ML-based network optimization. Algorithm training is considered one of the main obstacles to realizing AI-enabled wireless networks, which is usually very time-consuming. By contrast, LLM has shown impressive few-shot or even zero-shot learning capabilities in many fields[182]. In particular, LLMs can learn in context from few or zero network management and optimization examples and then generalize to incoming new tasks. By providing a handful of examples, the LLM agent can quickly learn the hidden patterns without any extra model training and fine-tuning, saving considerable time and effort for algorithm training in network management. In addition, such fast learning capability is critical to make rapid responses to network dynamics. This means that network optimization decisions can be efficiently adjusted based on traffic patterns, user types, and operator demands.